OpenStudy (anonymous):

identify the identity sec^2 [(pi/2) - x]- 1= cot^2 x please help!

OpenStudy (anonymous):

there it is...

OpenStudy (anonymous):

write that left side as \[\large \frac {1}{\cos^2( \frac {\pi}{2}-x)}-1\]

OpenStudy (anonymous):

then try using the sum/difference formulas for sin/cos...

OpenStudy (anonymous):

thanks!